OpenStudy (df001):

(1+sinΘ)(1-sinΘ)=cos^2Θ

OpenStudy (df001):

Help pls how do i make left side of the equation to equal to the right side

OpenStudy (df001):

hello

OpenStudy (johnweldon1993):

Just expand the left hand side and you'll see it in fact is true \[\large (1 + sin\theta)(1 - sin\theta)\] \[\large 1 - sin\theta + sin\theta - sin^2\theta\] \[\large 1 - sin^2\theta\] So we have \[\large 1 - sin^2\theta = cos^2\theta\] Is this true? What do we know from the trig identity \(\large sin^2\theta + cos^2\theta = 1\)

OpenStudy (df001):

ohh I see now!\

OpenStudy (df001):

so it is expandable!

OpenStudy (df001):

thanks

OpenStudy (johnweldon1993):

No problem!

OpenStudy (df001):

i thought i had to use identities lol

OpenStudy (usukidoll):

for these trig proofs you can use identities or expand... there are a lot of methods to try and make the left hand side = right hand side. It's just that some trig proofs are easier than others.